# CaloPointFlow II

### v1

@@ 1 Introduction
To tackle these challenges, there is an increasing effort to develop generative machine learning models as a more computationally efficient yet accurate method for calorimeter simulation. The objective is to develop models that can precisely reproduce the distribution of calorimeter responses while significantly reducing computational time and resources. Such an approach not only reduces the burden on the computational infrastructure but also has the potential to enable more intricate and precise analyses than those currently feasible.
To address these aspects, there is a growing field of research about deep learning based surrogate models [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 23, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] such as Generative Adversarial Networks (GANs) [15, 16, 17, 18, 19, 20, 21, 22, 23, 26, 27, 23, 36, 39, 40, 48, 49], Variational Autoencoders (VAEs) [21, 26, 33, 45], normalizing flows [24, 25, 31, 34, 35, 43, 45, 46], and diffusion models [28, 29, 37, 38, 41, 42, 47], that have been developed for detector simulations.
@@ 2 Datasets
In our research, we exclusively used the second and third dataset from the CaloChallenge [62]. Each shower in the dataset contains the incident energy of the incoming particle and a vector containing the voxel energies. The incident energy $E_{\rm inc}$ is log-uniformly distributed between $11\,\mathrm{GeV}$ and $11\,\mathrm{TeV}$. Each dataset consists of 200,000 showers initiated by electrons. These datasets are equally divided for training and evaluation purposes.
Datasets 2 [63] and Dataset 3 [64] are simulated using the same physical detector, which consists of concentric cylinders with 90 layers of absorber and sensitive (active) materials, specifically tungsten (W) and silicon (Si), respectively. Each sub-layer consists of $1.41\,\mathrm{mm}$ of W and $0.31\,\mathrm{mm}$ of Si, resulting in a total detector depth of $153\mathrm{1}]{mm}$. The detector’s inner radius is $801\,\mathrm{cm}$.
@@
The calorimeter geometry of Dataset 2 comprises 45 concentric cylindrical layers stacked along the direction of particle propagation ($z$). Each layer is further divided into 16 angular bins ($\alpha$) and nine radial bins ($r$), resulting in a total of $45\times 16\times 9=6480$ voxels in each shower.
Dataset 3 from the CaloChallenge features a higher granularity compared to Dataset 2. Each layer in Dataset 3 consists of 18 radial bins and 50 angular bins, resulting in a total of $45\times 50\times 18=40500$ voxels in each shower.
@@ 3 Model
Figure 1: A schematic view of the model. Part a) displays the training setup. The *Encoder* encodes the points into the latent representation $z$. The *LatentFlow* is optimized to learn $z$, while the *PointFlow* is optimized to learn the distribution of points $x_{i}$ conditioned on $z$. Part b) displays the sampling setup. First, the conditional variables $E_{\text{sum}}$ and $n_{\text{hits}}$ are generated. Then, the *LatentFlow* generates the latent representation $z$, which is used to generate the points with the *PointFlow*.
@@
A schematic view of the sampling process is shown in Figure 0(b). For each shower, first $E_{\textrm{sum}},n_{\textrm{hits}}$ are sampled by the CondFlow. A latent vector $z$ is sampled by the LatentFlow. At last, $n_{\textrm{hits}}$ points are generated by the PointFlow.
@@ 5 CDF-Dequantization
Transforming a uniform distribution into any distribution causes a noteworthy problem in the domain of statistics and data science. One solution is using the inverse transform $\Phi_{X}^{-1}(u)=\inf\{x|F_{X}(x)\geq u\}$. This transform maps $U\to X$, where $X$ is any distributions. A proof for the case of a continuous distribution is provided in Appendix B or can be found in the statistics literature.
@@
In summary, we have constructed a relatively simple transformation that can map discrete data to a normal distribution. In doing so, we have significantly eased the task of the normalizing flow, which is now left with only the responsibility to learn the correlations in the data and not the general shape. This enhances the efficiency and effectiveness of machine learning models dealing with discrete data. For a visual representation, see Figure 2
We provide the algorithms for both directions of the CDF-Dequantization below.
@@ 6 DeepSetFlow
The new flow architecture is able to capture the point-to-point correlations This approach shares similarities with previous work by Buhmann et al. [51] who constructed the EPiC-GAN, a model that applied DeepSets to create a permutation equivariant generative model. Mikuni et al. [54] also employed a similar technique of information aggregation within a diffusion model to learn point clouds.
In another instance, Käch et al. [53] developed a model that combines all information into a single node, referred to as the mean field. Their models main distinguishing factor from previous ones is the use of cross attention to update the information of the mean field.
@@ 8 Results
Figure 3: The distribution of the sum of all energies in the calorimeter divided by the energy of the incident electron for both datasets.
Figure 3 show the distribution of the sum of all energies in the calorimeter cells divided by the energy of the initial particle. A good agreement between Geant4 and CaloPointFlow II is concluded from this comparison.
@@
Figure 4: The distribution of cell energies for both datasets.
In Figure 4 we present the energy distributions of individual cells in the two different datasets. These figures show consistency within the bulk of the distributions across both datasets. However, one observation can be made regarding the tails of these distributions. In both cases, the CaloPointFlow II model shows a tendency to overshoot and undershoot in energy. This phenomenon is particularly pronounced in Dataset III, where the deviations from the expected values are more pronounced in the tail regions.
@@
Figure 5: The distribution of number of hits for both datasets.
The figures Figure 5 show the distribution of the number of hits in both data sets. The ”number of hits” is defined as the number of calorimeter cells that have an energy value greater than zero for a given shower. A key observation in both datasets is the generally accurate modeling of the overall distribution of hits. However, a closer examination of Figure 5 reveals a limitation in the models ability to reproduce the highest number of hits. This problem is most likely due to the multiple hits workaround. In scenarios where most detector cells are hit, the probability of mapping multiple points to a single cell increases significantly. This results in a discrepancy in the tail of the distribution, where the model fails to generate the extreme values observed in the real data. Conversely, for Dataset III this particular problem is not apparent. This dataset is characterized by a smaller proportion of total cells being hit. This observation suggests that the problem of accurately modeling the highest number of hits is less pronounced in more granular calorimeters.
@@
Figure 6: The marginal profiles of the average shower shapes in $z,\alpha$ and $r$ for both datasets.
Figure 6 show the marginal profiles of the average shower for the second and third data sets, respectively. The first plots in Figure 6 show the longitudinal shower profile. We observe minimal variation between the two distributions, indicating a strong agreement in the longitudinal profile. Moving on to the second plots in Figure 6, we examine the flat distribution of $\alpha$. That both distributions appear to be uniformly flat is an inherent property of our construction. Finally, the third plot in both figures examines the radial distribution, where we find that both distributions have similiar decreasing structure.
@@
Figure 7: The energy distribution in different layers of the detector for both datasets.
Figure 7 provide a detailed view of the distribution of total energy in different areas of the detector. Each figure consists of five plots, each showing the energy distribution in five successive layers of the detector. A consistent pattern observed in these figures is the accurate modeling of the energy distributions in the initial layers of the detector. However, there is a notable discrepancy in the later layers. In these layers, there is a significant absence of higher energy values in the model results.
@@
Figure 8: The Pearson correlation coefficients between all cells of the detector are displayed for both datasets.
Figure 8 show the Pearson correlation coefficients, which quantify the relationships between the energies of all cells in the detector. For visualization purposes, the three-dimensional structure of the detector is flattened, resulting in a recurring pattern within the displayed data. Each figure is divided into three distinct sections: the left section shows the energy distribution as simulated by Geant4, the middle section illustrates the results from the CaloPointFlow II model, and the right section highlights the differences between the two models. A key observation from these figures is the overall effective modeling of the correlation coefficients by the CaloPointFlow II model. This suggests that the model is able to capture the inter-cell energy relationships within the detector.
@@
Figure 9: Energy-weighted shower center distribution in all three dimensions for both datasets.
Figure 9 show the energy-weighted mean of the shower in Dataset II and III, respectively. The means $\mathbf{\mu}_{i}$ are calculated as
@@
where $\mathbf{c}^{i}_{j}$ are the 3-dimensional coordinate vector of the hit constituents of the shower and $e_{j}^{i}$ are the corresponding energies. The modeled data captures the general trend of the energy-weighted mean, but there are some discrepancies with the actual data. These differences may indicate limitations in the models ability to accurately simulate the nuanced distribution of energy within the showers.
@@
Figure 10: The eigenvalues of the covariance matrix weighted by energy are presented for both datasets.
@@ 9 Conclusion
In this study, we have introduced CaloPointFlow II, a refined generative model that significantly advances the simulation of calorimeter showers. The utilization of point clouds, owing to their inherent sparsity, offers a computationally more efficient approach compared to traditional fixed data structures like 3D images. Building upon the foundations of the original CaloPointFlow model, CaloPointFlow II incorporates innovative techniques such as a novel dequantization method, referred to as CDFDequantization, and a new normalizing flow architecture, named DeepSetFlow. These enhancements contribute to a more efficient and accurate modeling of particle interactions in calorimeters, enabling more efficient and accurate modeling of complex particle interactions.
The extensive evaluation of CaloPointFlow II using the challenging CaloChallenge datasets II and III has proven its efficacy. Our model not only exhibits superior performance in both low-level and high-level classifications compared to its predecessor but also showcases significant improvements in the accuracy of calorimeter shower simulations. The results demonstrate not only its capability to accurately simulate calorimeter showers but also a marked improvement in computational efficiency compared to traditional methods. The model exhibits an impressive balance between fidelity in the simulation and the computational demands, especially evident in the precise modeling of the spatial structure and energy distributions within calorimeter showers.

### v2

@@ 1 Introduction
To tackle these challenges, there is an increasing effort to develop generative machine learning models as a more computationally efficient yet accurate method for calorimeter simulation. The objective is to develop models that can precisely reproduce the distribution of calorimeter responses while significantly reducing computational time and resources. Such an approach not only reduces the burden on the computational infrastructure, but also has the potential to enable more intricate and precise analyses than those currently feasible.
To address these aspects, there is a growing field of research about deep learning based surrogate models [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 23, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50] such as Generative Adversarial Networks (GANs) [15, 16, 17, 18, 19, 20, 21, 22, 23, 26, 27, 23, 36, 39, 40, 48, 49], Variational Autoencoders (VAEs) [21, 26, 33, 45], normalizing flows [24, 25, 31, 34, 35, 43, 45, 46], and diffusion models [28, 29, 37, 38, 41, 42, 47, 50], that have been developed for detector simulations.
@@ 2 Datasets
In our research, we exclusively use the second and third dataset from the CaloChallenge [63]. Each shower in the dataset contains the incident energy of the incoming particle and a vector containing the voxel energies. The incident energy $E_{\rm inc}$ is log-uniformly distributed between $11\,\mathrm{GeV}$ and $11\,\mathrm{TeV}$. Each dataset consists of 200,000 showers initiated by electrons. These datasets are equally divided for training and evaluation purposes.
Datasets 2 [64] and Dataset 3 [65] are simulated using the same physical detector, which consists of concentric cylinders with 90 layers of absorber and sensitive (active) materials, specifically tungsten (W) and silicon (Si), respectively. Each sub-layer consists of $1.41\,\mathrm{mm}$ of W and $0.31\,\mathrm{mm}$ of Si, resulting in a total detector depth of $1531\,\mathrm{mm}$. The detector’s inner radius is $801\,\mathrm{cm}$.
@@
The calorimeter geometry of Dataset 2 comprises 45 concentric cylindrical layers stacked along the direction of particle propagation ($z$). Each layer is further divided into 16 angular bins ($\alpha$) and nine radial bins ($r$), resulting in a total of $45\times 16\times 9=6480$ voxels.
Dataset 3 from the CaloChallenge features a higher granularity compared to Dataset 2. Each layer in Dataset 3 consists of 18 radial bins and 50 angular bins, resulting in a total of $45\times 50\times 18=40500$ voxels.
@@ 3 Model
Figure 1: A schematic view of the model. Part a) displays the training setup. The Encoder encodes the points into the latent representation $z$. The LatentFlow is optimized to learn $z$, while the PointFlow is optimized to learn the distribution of points $x_{i}$ conditioned on $z$. Part b) displays the sampling setup. First, the conditional variables $E_{\text{sum}}$ and $n_{\text{hits}}$ are generated. Then, the LatentFlow generates the latent representation $z$, which is used to generate the points with the PointFlow.
rar
@@
A schematic view of the sampling process is shown in Figure 1(b). For each shower, first $E_{\textrm{sum}}andn_{\textrm{hits}}$ are sampled by the CondFlow. A latent vector $z$ is sampled by the LatentFlow. At last, $n_{\textrm{hits}}$ points are generated by the PointFlow.
@@ 5 CDF-Dequantization
Transforming a uniform distribution into any distribution causes a noteworthy problem in the domain of statistics and data science. One solution is using the inverse transform $\Phi_{X}^{-1}(u)=\inf\{x|F_{X}(x)\geq u\}$. This transform maps $U\to X$, where $X$ is any distribution. A proof for the case of a continuous distribution is provided in Appendix B or can be found in the statistics literature.
@@
In summary, we have constructed a relatively simple transformation that can map discrete data to a normal distribution. In doing so, we have significantly eased the task of the normalizing flow, which is now left with only the responsibility to learn the correlations in the data and not the general shape. This enhances the efficiency and effectiveness of machine learning models dealing with discrete data. For a visual representation, see Figure 2 We provide the algorithms for both directions of the CDF-Dequantization below.
@@ 6 DeepSetFlow
The new flow architecture is able to capture the point-to-point correlations. This approach shares similarities with previous work by Buhmann et al. [52] who constructed the EPiC-GAN, a model that applied DeepSets to create a permutation equivariant generative model. Mikuni et al. [55] also employed a similar technique of information aggregation within a diffusion model to learn point clouds.
In another instance, Käch et al. [54] developed a model that combines all information into a single node, referred to as the mean field. Their model’s main distinguishing factor from previous ones is the use of cross attention to update the information of the mean field.
@@ 8 Results
Figure 3: The distribution of the sum of all energies in the calorimeter divided by the energy of the incident electron (left) for Dataset II and (right) for Dataset III.
Figure 3 shows the distribution of the sum of all energies in the calorimeter cells divided by the energy of the initial particle. A good agreement between Geant4 and CaloPointFlow II is concluded from this comparison.
@@
Figure 4: The distribution of cell energies (left) for Dataset II and (right) for Dataset III.
In Figure 4 we present the energy distributions of individual cells in the two different datasets. These figures show consistency within the bulk of the distributions across both datasets. However, one observation can be made regarding the tails of these distributions. In both cases, the CaloPointFlow II model shows a tendency to overshoot or undershoot in energy. This phenomenon is particularly pronounced in Dataset III, where the deviations from the expected values are more pronounced in the tail regions.
@@
Figure 5: The distribution of number of hits (left) for Dataset II and (right) for Dataset III.
Figure 5 shows the distribution of the number of hits in both data sets. The ”number of hits” is defined as the number of calorimeter cells that have an energy value greater than zero for a given shower. A key observation in both datasets is the generally accurate modeling of the overall distribution of hits. However, a closer examination of Figure 5 reveals a limitation in the models ability to reproduce the highest number of hits. This problem is most likely due to the multiple hits workaround. In scenarios where most detector cells are hit, the probability of mapping multiple points to a single cell increases significantly. This results in a discrepancy in the tail of the distribution, where the model fails to generate the extreme values observed in the real data. Conversely, for Dataset III this particular problem is not apparent. This dataset is characterized by a smaller proportion of total cells being hit. This observation suggests that the problem of accurately modeling the highest number of hits is less pronounced in more granular calorimeters.
{subcaptionblock}
@@
Figure 6: The marginal profiles of the average shower shapes in $z,\alpha$ and $r$ (top) for Dataset II and (bottom) for Dataset III.
Figure 6 shows the marginal profiles of the average shower for the second and third data sets, respectively. Figure 6 (left) shows the longitudinal shower profile. We observe minimal variation between the two distributions, indicating a strong agreement in the longitudinal profile. In Figure 6 (middle) the flat distribution of $\alpha$ is visible. Both distributions appear to be uniformly flat, which is an inherent property of the construction. Finally, Figure 6 (right) both figures examines the radial distribution, where we find that both distributions have similar decreasing structure.
{subcaptionblock}
@@
Figure 7: The energy distribution in different layers of the detector (top) for Dataset II and (bottom) for Dataset III.
Figure 7 provides a detailed view of the distribution of the total energy in different areas of the detector. For each dataset, five plots are shown with the energy distribution in five successive layers of the detector. A consistent pattern observed in these figures is the accurate modeling of the energy distributions in the initial layers of the detector. However, there is a notable discrepancy in the rear layers. In these layers, there is a significant absence of higher energy values in the model results.
{subcaptionblock}
@@
Figure 8: The Pearson correlation coefficients between all cells of the detector are displayed (top) for Dataset II and (bottom) for Dataset III.
Figure 8 contains the Pearson correlation coefficients, which quantify the relationships between the energies of all cells in the detector. For visualization purposes, the three-dimensional structure of the detector is flattened, resulting in a recurring pattern within the displayed data. Figure 8 (left) shows the energy distribution as simulated by Geant4, Figure 8 (middle) illustrates the results from the CaloPointFlow II model, and Figure 8 (right) highlights the differences between the two models. A key observation is the overall effective modeling of the correlation coefficients by the CaloPointFlow II model. This suggests that the model is able to capture the inter-cell energy relationships within the detector.
{subcaptionblock}
@@
Figure 9: Energy-weighted shower center distribution in all three dimensions (top) for Dataset II and (bottom) for Dataset III.
Figure 9 shows the energy-weighted mean of the shower in Dataset II and III, respectively. The means $\mathbf{\mu}_{i}$ are calculated as
@@
where $\mathbf{c}^{i}_{j}$ are the 3-dimensional coordinate vector of the hit constituents of the shower and $e_{j}^{i}$ are the corresponding energies. The modeled showers capture the general trend of the energy-weighted mean, but there are some discrepancies with the actual data. These differences may indicate limitations in the model’s ability to accurately simulate the nuanced distribution of energy within the showers.
{subcaptionblock}
@@
Figure 10: The eigenvalues of the covariance matrix weighted by energy are presented (top) for Dataset II and (bottom) for Dataset III.
@@ 9 Conclusion
In this study, we have introduced CaloPointFlow II, a refined generative model that significantly advances the simulation of calorimeter showers. The utilization of point clouds, owing to their inherent sparsity, offers a computationally more efficient approach compared to traditional fixed data structures like 3D images. Building upon the foundations of the original CaloPointFlow model, CaloPointFlow II incorporates innovative techniques such as a novel dequantization method, referred to as CDF-Dequantization, and a new normalizing flow architecture, named DeepSetFlow. These enhancements contribute to a more efficient and accurate modeling of particle interactions in calorimeters, enabling more efficient and accurate modeling of complex particle interactions.
The extensive evaluation of CaloPointFlow II using the challenging CaloChallenge datasets II and III has proven its efficacy. Our model not only exhibits superior performance in both low-level and high-level classifications compared to its predecessor but also showcases significant improvements in the accuracy of calorimeter shower simulations. The results demonstrate not only its capability to accurately simulate calorimeter showers, but also a marked improvement in computational efficiency compared to traditional methods. The model exhibits a balance between fidelity in the simulation and the computational demands, especially evident in the precise modeling of the spatial structure and energy distributions within calorimeter showers.